HST Status Mar 07, 2011

Skip to first unread message

Dr. Paul Scowen

Mar 23, 2011, 12:27:53 PM3/23/11

Earlier reports Convert Day of Year

The Hubble Space Telescope

Unlocking the Secrets of the Universe

Daily Report #5349

Period Covered:
07:00 pm March 6, 2011 - 06:59 pm March 7, 2011
(DOY 66/0000z - 66/2359z)

Flight Operations Summary

Significant Spacecraft Anomalies: (The following are preliminary reports
of potential non-nominal performance that will be investigated.)

12744 - SCI: REAcq(2,1,1) at 04:34:51 failed with scan step limit exceeded on FGS 2 five times at
04:37:30, 04:38:44, 04:39:58, 04:41:21 and 06:16:29.@066/0437z
Observations affected: WFC3 6 - 7, proposal ID#12356, ACS 8, proposal ID# 12397.

19034-2 - Install ACS FSW CS 5.16, ASIC3.6 (JV0035) into EEPROM@66/1746z


FGS GSAcq		5		5
FGS REAcq		10		9
OBAD with Maneuver	5		5

The COS FUV segment B nominal operating HV preset was successfully updated in Executive RAM and 
copied to EEPROM. Additionally the maximum FUV HV FSW limits which were already in Exec RAM were 
copied to EEPROM. Segment B nominal operating voltage was increased from 167 to 175 counts (delta 
~125.6 volts). 

Ops Request 19034 was successfully completed at ~066/21:00 to load the latest release of the ACS CS 
FSW 5.16 (JV0035) into EEPROM. ACS CS FSW 5.16 will be activated via stored commanding tomorrow at 
~067/1900 when ACS is recovered from Safe mode following the installation of NSSC-1 BQ9.2

Program Principal Investigator Program Title
Science Observations Scheduled
11582 Andrew Blain, California Institute of Technology The spatial distribution of radiation in the complex ISM of distant ultraluminous galaxies
12161 David Ardila, Jet Propulsion Laboratory Accretion in Close Pre-Main-Sequence Binaries
12278 Thomas Ayres, University of Colorado at Boulder Advanced Spectral Library Project: Cool Stars
12286 Hao-Jing Yan, The Ohio State University Hubble Infrared Pure Parallel Imaging Extragalactic Survey {HIPPIES}
12376 Vinay Kashyap, Smithsonian Institution Astrophysical Observatory The Spinning Corona of FK Comae
Calibration Observations Scheduled
12342 Tiffany Borders, Space Telescope Science Institute WFC3 UVIS CCD Daily Monitor
12349 Michael Dulude, Space Telescope Science Institute IR Dark Monitor
12356 Norbert Pirzkal, Space Telescope Science Institute IR Grism Wavelength Calibration Stability Test
12397 David Golimowski, Space Telescope Science Institute CCD Daily Monitor {Part 2}
12400 Michael Wolfe, Space Telescope Science Institute Dark Monitor Part 1
12402 Michael Wolfe, Space Telescope Science Institute CCD Bias and Read Noise Monitor-Part 1
12420 Wei Zheng, The Johns Hopkins University NUV Detector Dark Monitor

Reply all
Reply to author
0 new messages